OpenStudy (anonymous):

integrate this please. int dx / 1+cosx

OpenStudy (anonymous):

\[\int\limits dx / 1 + \cos x\]

OpenStudy (anonymous):

let u=tan(x/2)

OpenStudy (anonymous):

where did tan (x/2) came from?

OpenStudy (anonymous):

This come from the right angle triangle but you can times the top and the bottom by (1-cosx) <-- this will probably be the easiest way

OpenStudy (anonymous):

right. yes. i get it now. thank you very much. i will answer it first, ok? :D

OpenStudy (anonymous):

sure no worries :)

OpenStudy (anonymous):

i got it. thank you! :D